NAOMI/OASIS job list
NAOMI technical information · WHT · Astronomy · ING

Version: 2006 Dec 21

NAOMI/OASIS commissioning / enhancements jobs are summarised below under 8 headings:

Under each heading, the goals are listed in rough priority order, with some indication of which tasks are crucial, and which are to be attempted as time permits. Crucial goals are listed within each section before the line-break.
Further OASIS-specific commissioning tasks are included in CRAL's 13/6/03 plan for the Jul 2003 commissioning.

Most of the tasks below relate to one of the following critical issues ultimately affecting delivered image quality or observing overheads for NAOMI, OASIS, OSCA and INGRID:

This list does not include work related to commissioning of GLAS, and probably omits a number of outstanding INGRID and OSCA tasks (see lists maintained by instrument specialists).

Initials indicate who's doing what: Chris Benn (CRB), Craige Bevil (CB), Nigel Dipper (NAD), Stephen Goodsell (SJG), Tom Gregory (TG), Olivier Martin (OM), Richard Myers (RM), Sam Rix (SR)

NAOMI astronomy team ('astros' below) = CRB, SRIX, WJI
NAOMI engineering team = CB, MFB, OM, MvdH, PJ, SGR, SMT, SPICO, TG.

On-sky PSF performance tests


Useful lists of test targets can be found on the AO technical page.

  1. Characterise performance with new TT system. Does it meet GLAS requirements?
  2. Characterise performance (FWHM, Strehl, in-slit energy) in both IR and optical, as a function of guide-star R mag, separation between guide star and target (out to at least 1 arcmin), wavelength, natural seeing and lenslet/readout mode. It's not necessary to explore the whole parameter space on-sky, just enough to pin down the grid of model predictions. Compare performance on faint stars with and without moon. Check on-sky for Airy rings, FP has seen in JHK with other systems e.g. Calar Alto, CFHT (see images in Rigaut et al). In the optical, determine how much of correction seen at short wavelengths is tip-tilt. Use dichroic 2 (most-used, allows plenty of light through to the wavefront sensor).
  3. Test fast-guiding/static-cleanup mode 10 with doublet. (Easiest to lock static corrections with 8x8 mode, then fast guide with doublet - RM 7/01.) This mode will be very useful with OASIS - ~ half the FWHM reduction at optical wavelengths appears (7+8/04) due to fast tip/tilt - why? TT appears to be taken out well with DM-frozen mode 1, but not with mode 10 - problem is centroid errors due to speckle? (RWW)
  4. Investigate why we are not routinely reaching the diffraction limit in good seeing, with bright guide stars. In particular, why do modes 2, 3 not improve performance over mode 1? Check no large offsets for mode 2 (it's very sensitive to this, because quad-cell). Use new naomi reconstructor. Compare Gao's new modes 11 and 12 (8*8 binned to 4*4 or 2*2) with mode 1, preferably on a field containg several stars. Check what gains are loaded on changing modes.
  5. Test 4*4 mode on sky, demonstrate that it delivers an improvement in performance. New WFS modes 13 14 15 needed first.
  6. Sequence of open-loop / closed-loop images while guiding on bright star, to use as PR movie. Try a 0.1-arcsec binary.
  7. Check optimum gain for each mode.

  8. In ~ 1-arcsec seeing, record WFS data on bright star modes 1 and 4, so we can compute loss of performance due to light falling outside central pixels. Done 4/02? Also, if time, record PSFs in modes 2,5, to see whether mode 5 does give any gain.
  9. Measure PSF at 10 arcsec radius from a guide star, then again for a guide star ~ 1, 10, 30 deg away, to see how degradation changes with position on the sky.
  10. Check that NAOMI satisfies optical-preformance requirements for OASIS e.g. V = 10 guide star in median seeing delivers FWHM 0.4 arcsec in median seeing; V = 13 at 0.9 micron delivers 0.35 arcsec. ~ OK, but needs confirming.
  11. Test bias (gradient) correction on WFS CCD, suggested by RM. How important is this?
  12. Check visibility of diff spikes due to secondary vanes (confirm they rotate with selected sky pa).

Other on-sky work

  1. Observe TAC-allocated science programmes in service mode. The applicants provide observing details.
  2. OSCA tests, particularly throughput measurement and suppression performance, new Lyot stop with cross at correct angle (one of 3) .
  3. Repeat of throughput tests, with new derotator optics, including repeat of throughput to WFS with different WFS exp times, important for GLAS.
  4. Observation of standard through OASIS, to get zeropoint.
  5. Measure throughput to WFS and INGRID, and INGRID background in J, H and Ks (all done, but worth redoing). SR 2/05 notes discrepancy between INGRID/Cass and INGRID/Nas numbers.
  6. Try SR's test targets.

  7. Check vignetting when the 20" OSCA field is offset to be centred on a quadrant (raised by SE 3/04).
  8. Check that AOWFAVSE header is recording WFS seeing correctly.
  9. Measure sky background on WFS in full moon. Check that it varies linearly with requested exposure time. Done, according to IKS. Can do with lit dome?
  10. Check scale "/pixel on MIT CCD.
  11. Measure throughput of OASIS to MIT CCD with each dichroic.
  12. PR images (ideally 3-colour) + proper reduction (and 3-colour rendering).
  13. Compare systematically the WFS, RoboDIMM and INGRID open-loop seeing measurements. 9/03: error in seeing calculation corrected (NOM).
  14. Check that dither offsets are in TCS headers, as well as the absolute positions [CB, PIMS 230].
  15. Determine for what modes the seeing monitor value gives reliable measurements.
  16. Check that seeing monitor gives plausible measurements during very bad and very good seeing, as well as during typical seeing. Check the display of seeing vs time on CB's gui.
  17. Test remote operation from SLO (see DA/CB emails 2/7/01) [CRB]. New ds9-based display tool needed? NAD advises setting decimate ~ 0, aim for one frame per sec. line to mountain is 2Mbuad?
  18. Check that guide stars anywhere in the 3-arcmin field can be used. Check that guide star is well-focussed on WFS at edge of field (automatic compensation of focus by NAOMI), and measure intensity of guide star on CCD to assess vignetting. This is OASIS requirement 5.
  19. Check no vignetting at INGRID, as function of star position (use the moveable pinhole in new NCU). [PIMS 280]
  20. Open-loop WFS centroid data immediately before an observation, required for Juan Carlos Guerra (TNG).
  21. Check that automatic log and FITS headers now include all reqd items, especially OSCA, lenslet, pupil stop. 10/03 FITS OK for OASIS, but log has no entries (req SGR 23/4/03). Use 11-char 'Filters etc' for OASIS short-form config name (= filter name, 5 chars) + 4 chars for pixel scale in arcsec (req SGR 23/4/03)
  22. Identify origin of left-right up-down cross-shaped (but unbeaded) diffraction pattern on INGRID, noted by SE for bright stars.
  23. Measure const of proportionality between defocus of tel and on WFS.
  24. Commission WFS ADC, maybe.

Off-sky PSF performance tests by astro team

  1. Measure up FWHM, 50%-eed and Strehl from commissioning and science images. Test independent prog (e.g. Stefan Hippler's IDL routine) for measuring strehl? How much does measured strehl depend on bgr slopes? How did Durham measure bgr on the 40*40 commissioning images?
  2. Controlled test of effect of humidity
  3. Compare the above with predictions from Ilona's model (which takes into account segmented mirror). Identify factors limiting performance.
  4. Monitor stability of optical-wavelength PSF on AG5 (following mirror event 2003 Oct 8). Temp comensation off.
  5. Monitor alignment (regular measurements) using grid of pinholes in front of INGRID.
  6. Further development of Ilona's performance model .
  7. Check WFS read noise and gain.
  8. Investigate whether humidity variations are responsible for frequent losses of DM shape 9/04. Can the airco control humidity? 10/04 surfaces dewed up, RGT considering domestic de-humidifier (18/10/04).
    Example DM change of shape in file.

  9. Investigate and ameliorate cause of temperature spikes near DM when cryostats filled.
  10. Check DM temperature compensation. Probably completely wrong, accumulated errors, needs redoing. Summer 2004, switched off. Does it get loaded on namirestart? Or is there a 10-min pause? If the latter, need to get it loaded on restart. Note from Paul Clark 1/01: [The gauge factor (ie. the gain) of the strain gauges changes with temperature so the model should be a curve (?) fit not just a straight line fit. Need to model temperature against: the actuator length, the strain gauge resistance and the gauge factor (the gain of the strain gauge system). Perhaps we should get a modeller to have a go at this from a theoretical point of view and then check it against reality?] Note from Richard Myers 1/01: [Test load-balanced strain-gause DFM feedback code. Will improve temporal response. This is still under development but can be gotten ready for a particular date.]
    Take new (better sampled) strain gauge response data. This is to build a MATLAB/Simulink model of the control system (Martin Fisher has offered to help us with this). Such a model will eneable us to more easily quantify the performance/wavelength range benefits of putting in fancier control algorithms (eg. Smith Predictors) and/or porting the control system to a new generation of DSPs (probably C67s).]
  11. Decide what kind of sky baffle is needed. 2", 4", none, or configurable? Will be an advantage when using the L3 WFS?
  12. Identify what is optimum MF for given window size in each of optical, J H and K.
  13. Investigate why mode 8 doesn't deliver better correction than mode 1 even for faint stars V > 14. Check that counts/cell in mode 1 from lamp are consistent with those in mode 1. Put sine wave on FSM, compare correction delivered by modes 1 and 8. Use TT injector in NCU?
  14. Investigate why mode 9 causes loss of guiding. Is direction in which star zooms off always the same? Check that amount of light in mode9 spots looks OK compared with other modes.
  15. Compare PSFs through different dichroics, investigate how much simplexing is needed to tweak up the PSF between dichroics. Try adding Zernike terms to simplex obtained with one dichroic, to optimise PSF on another dichroic. Try subtracting flat for one dichroic from flat from another, analyse in terms of Zernikes (these 2 suggestions from AJL). Probably no high-order terms from dichroics themsleves, but may be wavelength-dependent terms in optics downstream, acc to TG.
  16. Check dark counts and readnoise on WFS CCD, combine with measured sky and star counts to confirm S:N expctations [CRB]. What happens to dark counts on binning? Does the centroiding sware understand this?
  17. Look at extent to which simplex in one wavelength can be adjusted for another wavelength by adding low-order aberrations (after making sure no lambda ambiguities have been eliminated). (AJL email 23/4/03).
  18. Check with flat in front of DM, pinhole illumination,
    (1) can still focus on OASIS? what shift?
    (2) degradation acceptable for non-AO work?
    (3) lateral shift OK (2 nuts top-left, bottom-right to define position).
    Could be Ag-coated, ~ 1k euros (TG 3/05)
  19. OSCA - what contrast ratio can get as function of radius, in J, H
  20. Explore image degradation near edge of INGRID chip (degrades fwhm = 0.23 to 0.39 arcsec as move from radius 10 to 15 arcsec from centre of chip (factor 5 times too big to be due to known tilt of chip, according to mfb's analysis). Use moveable pinhole.
    Roy's report from 2/10/03: Done in x-direction. A series of images shows no obvious degradation. The images are named r627000 to 012. The movable pinhole does not move far enough when it is in naomi-simplex position. It reaches its limit at pixel 700, so I switched to the oasis-continuum (same 2um pinhole) to do the rest of the series (images 007 - 012). Steps are 1 unit on the slide which corresponds to about 115 pixels on ingrid.
  21. Check that the DM can be simplexed such that the pinhole image is diff-limited in J (as it was in Durham in R). Investigate effect of simplexing with 4 central elements obscured [new mask req 1/03]. Hartmann screens on OAP, to improve focus? [TG] Measure scale on TEK uing FSM to offset pinhole. Check how much strehl reduced due to light lost into spikes due to DM diffr pattern [SE]. Why does change between 2-mic and 5-mic pinholes give repeatable focus shift of 3 cm at TEK?

  22. Review segmented DM vs cont face-sheet, by plugging into Ilona's model the rms wavefront errors predicted for each.
  23. Why is PSF H/Fe2 different from H/clear?
  24. Devise recipe for deciding, on basis of guide-star mag and seeing, which mode to use.
  25. Look at differences between successive mirror flats, e.g. that between laser and wl flats appeares to have delta piston quantised in units of 800 DAC??
  26. Write software to add Zernike terms to flats, generate new flat file.
  27. Look at differences between successive offset files, appear to be dominated by tip/tilt wander. 25- and 2.5-msec offset files are very similar. Need to measure strehl on pinhole, closed loop, as function of peak heights on WFS. Explore use of mean offsets file (rather than measuring each time).
  28. Laser flat, close loop with no offsets, recon of, counts rise 20k -> 30k (pres because takes out residual tilts), open loop with 'load flat on open' off, counts are the same, save DM, now says it's 'loaded' but counts are the same, now load explicitly from new, and counts drop back to 20k. Why? DAC values are unchanged, although spots move on WFS.
  29. Test recovery of PSF by defocussing WFS then science camera??
  30. Test sor8 reconstructor
  31. Commission modal control rather than zonal? Most straightforward for 4*4. Probably not much gain from this, according to Ilona's models.
  32. Where does simplex (maximising sum of I^2) try to position peak? Looks as though it strives to spread the counts evenly over 4 central pixels. Simulate.
  33. Investigate PSF-measurement / data-analysis strategies. In particular, can we recover the PSF from the mean and variance of the spot positions, i.e. 2 * 72 (2 * 52 in practice) numbers? Should we record these in the FITS headers? PSF pinhole + measure of r0 is sufficient? [SE].
  34. Measure quality of psf as function of position in the field, using moveable pinhole.
  35. Conjugate to turbulent layer for *2 extra sky coverage with same strehl? Dicsussed 6/3/01.

Other off-sky tests by astro team

  1. Get CB's DM monitor working, check DM not moving in mode 10
  2. Identify test star fields all RA
  3. Mode 10 loop speed
  4. Check ADC formulae
  5. Characterise flickering of NCU lamps, for investigation
  6. Sort out how to maintain OSCA/NAOMI/INGRID alignment.
  7. Test reliability of WFS mechs esp foreoptics
  8. Review strategy for reducing emissivity in K (from derotator mirror and windows and from INGRID fore-optics) [TG/AZ]. Comparisons of emissivity through the two sets of foreoptics need to be done with INGRID out of GRACE. Lack of AR coating can't affect emissivity of foreoptics (TG) - more likely that INGRID is seeing the foreoptics casing.
  9. Commission OASIS ADC, when installed.
  10. Check repeatability of positioning one of dichroics in the holder (from position of pinhole image). Check repeatability when changing dichroics. Do after cone/v/flat mount installed (not yet ordered 21/10/04 acc to mfb).
  11. Discuss solutions to 100% vignetting by OSCA of beam to OASIS, when OSCA raised.
  12. Check counts/sec on OASIS/HR from a dome lamp.

  13. Measure WFS gains
  14. Establish how to reboot fisba from cont room
  15. Monitor sky baffle for slippage
  16. Measure OSCA mask opacities in Z J K using NCU pinhole (Sebastian already did H, see his email 23/2/04).
  17. Consider cooling (a) INGRID foreoptics, (b) whole enclosure, especially OSCA surfaces, to reduce emissivity and improve DM stability. Predict effect on emissivity of cooling INGRID foreoptics to 0 deg C [MFB?].
  18. Identify which INGRID quadrants correspond to which quadrants on simplex gui.
  19. Measure MIT fringing (W lamp + narrow-band red filter). NB final test needs to be done with OASIS. SMT has already done for this chip? Peletier ITP data poor due to 5% fringing with MR516 (9/8/06 email from wji)
  20. Check CTE of MIT.
  21. Get dichroic curves on web. MFB will spot check transmission curves at 22 deg vs manufactured curves (agreed 6/02).
  22. Train in use of XOASIS [crb, wji].
  23. Test Savart-plate Shearing-interferometer view of DM (analogue view). Accurate positioning required. Analytical software probably not now required.
  24. Establish safe procedure for changing dichroics.
  25. Determine what objects make suitable test targets (globular clusters, binary stars etc.), put together a catalogue.
  26. Move OASIS web site to ING [GA,CRB]. 4/04 still being updated by Lyon.
  27. Measure read-noise on WFS at regular intervals.
  28. When air cond installed in GRACE, check no degradation of PSF due to vibration.
  29. Explore possibility of using laser in IR to halve number of fringes, reduce sensitivity to vibration. [TG?]
  30. Some sdsu modes may have reset times wrong (AJL 1/03), check linearity [astros, with ATC for follow-up].
  31. Try pistoning a segment from black thru white to black, check it moves by 6300/4 A ~ 200 DAC.
  32. Check autowhite loading flat afterwards (20/6/03).
  33. Experiment with deconvolution software.
  34. Decide which INGRID pupil stop should be loaded automatically with each INGRID filter (NAOMI and OSCA), by checking intensity and background for each filter/stop combination. Stop sizes are in Almu's email 25/4/02 [CRB/RO/SE]. Might be better to do this after any improvement of alignment of secondary and DM central 4 elements. Observations made 020826. Need pupil images of star through each pupil. Pupil wheel full, one stop must go? OSCA stops -> 3.4-m primary.
  35. Check repeatability of positioning table.
  36. Observing overheads calculator.
  37. Explore changes in mirror properties (i.e. how long does a simplex last) with time and temperature, with table up, with table down, when mirror idle, when mirror being exercised (in response to turbulence generator).
  38. Determine best strategy for flattening the mirror (a) slowly, (b) in a hurry, from different starting conditions (e.g. from a previous simplex, or from mid-range). Is the difference flat between the white-light and simplex flats fairly constant, i.e. are non-common-path errors stable enough to take out without lookup table? [CRB] This item can wait until after tests with the Savart plate?
  39. Train up on how to fix FISBA when the fringes vanish [TG].
  40. Determine under what circumstances AutoConfig sets WFS integration time to zero (RM email 3/7/02) [CRB].
  41. WFS ADC daytime tests: check position angle by looking at image on doublet (GG495 out). Transfer ADC positioning code to our machines? [AJL?] Test ADC on sky. Not essential for NAOMI. Important at visible wavelengths, but OASIS has own ADC, but if this fails, NAOMI ADC needed for WFS (at OASIS science still recoverable even if dispersion uncorected) [TG/CRB, July for off-sky test with doublet]. Contact AJL after 2/03 for details of test results (his email 16/1/03). **
  42. Investigate ghosts in INGRID, using moveable pinhole image to sample ghosting as function of position of bright star. Ghosting may depend on number of filters (1 or 2) in beam. [CRB]
  43. Cover lights on laser. Measure light pollution on CCD with and without LEDs covered. [PIMS 271]
  44. Check why iframediff and ilistdark sometimes fail on taurus [CRB].
  45. Display NAOMI info and PR images above laser printer in WHT [CRB, done].
  46. Complete analysis of throughput to WFS:
    1. Does empirical starlight throughput match that expected from throughput measurements made on individual components e.g. scan of dichroic, RG630 filter?
    2. Evaluate gains possible through recoating etc. [PIMS 273]
    3. What is thruput of Nas flat at 45 deg incidence?


Low priority:
  1. Remove outer 16 segments of simplex to speed up? Not much gain.

Off-sky software work

Most of these are also included in the PIMS database.

NAVIS

  1. 1-keyboard control (of 2 or 3 screens)

  2. Get IDL updated to 6.0 (req by R McDermid 3/05).
  3. .postcshrc needs cleaning up (Gary's email of 14/2/03)
  4. lpr command no longer recognised on navis, as of 14/2/03.
  5. New screens for AO - NB the 19-in sun displays are fuzzier than the 20-in Dell displays.

RTCS / Top Gui

  1. Indicator showing when loop is *really* closed, i.e. when corrections are being sent.
  2. Get reconstructor loaded on startup [RM].
  3. Install settable S:N limits on DM segments, and optional fallback to mirror-flat positions when S:N low.
  4. TopGui crashes if > 20 unique configurations saved on scripts page.
  5. C40comms crashes appear to be caused by NaomiAutoguider script (RO email 14/10/03), due to it bypassing the EPM layer and retrieving centoids directly from C40s. It should be using the EPM. Is there a bug in the script (some resource not properly freed up) that can explain the crashes?
  6. Int time > 1 sec for WFS (useful for non-AO guiding, not AO) req 5/05 (DR ??).
  7. New mode on WFS 8*8-> 2*2, 4*4 requested from Gao 18/6/03. Same needed for 4*4 lenslets (no point testing 4*4 until these are ready), requested 7/03. Who at ING? Code = 200 lines assembler. SJG visiting ATC Nov 03 for training.
  8. Double SG-loop bandwidth (curr 3 kHz) to reach predicted performance (straightforward software mod). Test running at 5 kHz 9/03 [RM/SJG]. Further off-sky work needed when RM out Nov 03.
  9. Test feed-forward operation of RTCS to increase effective bandwidth by predicting next mirror movement on basis last one [SJG visiting Durham to do this?].
  10. Put jittery elements on strain-gauge pass-thru mode during laser flattening only (currently they are on permanent pass-thru) [SJG/RM, PIMS 239].
  11. Display some measure of S:N on centroids, and define threshold which ensure that guiding stops if S:N drops below a given value. What happens at the moment if S:N is poor? Display of sum of counts in WFS cell selected by cursor. Display of sum of cells. [SJG/NAD/RM, PIMS 238]
  12. Averaged readout of WFS for e.g. on-sky focussing.
  13. Can get gain on WFS speed by not reading pixels outside pupil?
  14. Put NAOMI software into CVS [CB, SJG, Durham].
  15. work.gs is missing, so 3D mirror plot prog won't run [PIMS 247] [CRB to send name of prog to NAD]. WhiteLightPlot also fails, dislin package missing (see RO email 27/5/04)
  16. Zernike added at command line doesn't reflect in boxes on gui [NAD, PIMS 249].
  17. AutoOffsets to default to modes g1, 4*4 only, mode9 only [SJG, PIMS 254].
  18. Check AutoTest runs with one mode only [PIMS 233, fixed 3/03].
  19. aocontrol1 training for ING staff [RM].

  20. Indication on top gui of when DM and FSM are powered on (req Olivier 18/11/05).
  21. Upgrade python to v1.5.2 (CB email 24/7/03).
  22. Check if topgui really crashes on DASRESET [CRB, PIMS 245].
  23. Put Auto scripts under TopGui, some progress made 4/03 [CB/RM?, PIMS 256]
  24. Investigate why EPM delivers bursts of high-seeing (> 6-arcsec) values to Craige's seeing vs time plot e.g. 13/11/02 [NAD/CB].
  25. WFStoFITS fails if exp time > 1000 msec (emailed NAD Nov 13).
  26. Implement focus offload to TCS (not done at 7/02, acc to FJG).
  27. Averaged readout of WFS for e.g. on-sky focussing.
  28. Can get gain on WFS speed by not reading pixels outside pupil?
  29. Put NAOMI software into CVS [CB, SJG, Durham].
  30. work.gs is missing, so 3D mirror plot prog won't run [PIMS 247] [CRB to send name of prog to NAD]. WhiteLightPlot also fails, dislin package missing (see RO email 27/5/04)
  31. Zernike added at command line doesn't reflect in boxes on gui [NAD, PIMS 249].
  32. Option to set WFS exposure time manually, so can set with better precision. [CB to investigate restrictions, CRB to decide what new exposure times would be most useful, PIMS 232].
  33. AutoOffsets to default to modes g1, 4*4 only, mode9 only [SJG, PIMS 254].
  34. Check AutoTest runs with one mode only [PIMS 233, fixed 3/03].
  35. Commission 2-CCD mode (may be superfluous if get L3 CCD, see RM email 3/4/02).
  36. Vertical scale on the .py scripts which plot e.g. differential mirror flats [SJG/NAD, PIMS 258].
  37. FSM power spectrum button creates an empty graph [SJG/NAD, PIMS 260]. NAD 9/03 couldn't reproduce it, check.
  38. Should be able to drive WFS focus via gui. More than 1 week's work, according to NAD, no time to do it 4/03 [SJG/NAD, PIMS 262]

Configuration-control (excluding topgui)

  1. Light-path MIMIC
  2. OASIS gui:
    1. Stop button for irises and NCU
    2. Lamps switch off automatically after 1 hour
    3. Lamps switch off when NCU moved to new position
    4. Iris settable as ND, not encoder position
    5. Iris settings in database, looked up by gui
    6. Filter settings in database: F = for AG7 on wlflat DM; A = OASIS imaging MR661; B9F0 = offset from A for each filter (req sgr 16/6/05)
    7. Dichroic parks when shutdown selected on oasis gui
    8. OASIS flat IN/OUT -> OASIS/INGRID


  3. Overall control gui for NAOMI/OASIS/OSCA/INGRID/GLAS (combining light-path mimic and control)
  4. Go and cancel buttons on the oasis config page
  5. Add OASIS dichroic name to observing log, e.g. under 'Grating'
  6. Add OSCA
  7. Beam-steerer values real not integer, req sgr 10/1/05
  8. OASIS gui: 'imaging' not 'imagery' (req 10/03); NCU slide box sometimes says 'select' instead of naming the position (req 10/03); adc not adcds (req 9/04)

SG monitor

  1. Modify to display tilts rather than actuator positions.
  2. Check it doesn't crash C40s (see RM email 3/9/03).

  3. More labelling on the display.

Laser-flattening

  1. Make fisba software hunt for elements which have strayed a long way off. At present, if the fisba can't detect fringes on an element, it has to be tweaked in by hand. Fisba should do a spiral search in tip-tilt. [CB, PIMS 228/207]

  2. Investigate why fisba falls over so often [SJG, PIMS 203].

Simplexing

  1. Better PSF-quality monitor on display, e.g. plot of integrated flux vs radius from centroid [disc with Jure, 22/2/05].
  2. Should calculate strehl and other PSF metrics, as well as merit function, even if only every few sec.
  3. Should read out MF even when not simplexing.
  4. Need to see on simplex display when iterations get to point of moving segments by < 1 DAC (typically after ~ 2000 iterations acc to RO 30/10/03).

  5. Reversal of x button on simplex gui so that direction matches display [discused with Jure 22/2/05].
  6. When simplex finishes, should save mirror flat automatically [CB, PIMS 229].
  7. Get INGRID quadrants on the SIMPLEX gui in correct orientation [CB, PIMS 231].
  8. Start value of max counts in simplex is always higher than anything on frame, so it never changes. Why? (RM comments that this may be calculated by the sbig server.) [CB?, PIMS 234]

Target acquisition

  1. Craige's dither script to operate at non-zero sky pa, check whether transformation of WFS probe coords is around WFS zero point or optical axis, which is correct? (CB email 24/7/03)
  2. Craige's dither script should accept return as well as 'y'. Documentation needs to record sky pa limitation, and warning that can't use ampersand. Should note that for -multi, needs only n-1 x,y pairs. Should note that works for offsets in x,y not RA, Dec. Why do the offsets sometimes need to be specified, sometimes 0,0? [CB, PIMS 88]

  3. Add osca?

Data acquisition

  1. Get average seeing written into FITS headers [CB, PIMS 224, see Nigel's email 3/9/02, req confirmed to CB 10/03].
  2. When images read out, they should automatically display, start iraf imexam, provide statistics on the brightest image in the field, start up aoperf to measure strehl.
  3. Ensure that OASIS FITS headers are written at start of integration, e.g. currently OAFILT is written as 'busy' if changed immediately exposure terminates. Raised by RO/SGR 24/5/05.

XOASIS and exp-time calculator

  1. Version numbering
  2. See long list of bugs in Sam's Jan 2006 document
  3. Transfer OASIS exp-time calculator from Lyon to ING, needs NAG-substitute routines [left with RG, 5/05, now with Don 12/06].



Other off-sky engineering work

The actions below are grouped under:

GLAS-related

  1. Cover any LEDs in GLAS with something which *really* blocks photons, before integration with the rest of the AO system.

NAOMI optics

  1. Block light leaks from dichroic changer and other components containing LEDs (see measurements made 1/06), and make sure all LEDs in GLAS blocked before integration with NAOMI. Buy night-vision goggles ~ 250 e (dark-adapted eye insens to wavelengths > 6400 A, acc to ms 23/1/05).
  2. Turbulence generator for winter 04/05: hot wire or rotating disk? LCDs probably not useful because monochromatic, problems with diffraction. (But TG exploring poss of using cheap video display LCD which could be fed with comp-generated turbulence.) Rotating disk probably best to start with, mounted above NAOMI simplex pinhole.
  3. 4-arcsec sky baffle for testing in 04A. Ideally 2.5-arcsec and 4-arcsec sky baffles needed (the latter can be used when closing the loop on planets or galaxies with diffuse cores) [TG, PIMS 270].
  4. Left side of WFS box (as seen when standing at CCD end) needs new cover, to prevent dust getting in (and on encoder tracks), with easy-to-open inspection hatch, and sealed cable exit. Req 2/06.
  5. Position sensor for fisba shutter, info fed to lpath mimic. RGT 22/3/06 notes no effort for this until post-GLAS.
  6. ND = 1, 2 filters (absorption) for WFS (for bright stars with high-order AO, also for m <~ 10 stars for TT). TG investigating 8/3/04. To be installed 8/11/04 acc RGT schedule.
  7. Second small fibre feed in NCU, for K transmission [TG/MFB].

  8. More secure fixing for sky baffle, keeps slipping (and thus vignetting)
  9. Savart-plate interferometer Felix, manual operation by 4/04. [TG].
  10. Internal baffle in NCU [TG].
  11. OSCA self-vignetting means need thinner mount for which optic?
  12. Ag coat Nas flat in ING tank [TG/JR, PIMS 268].
  13. Ag coat 3 mirrors in derotator, here (or contract out, hand carry to Denton?), and replace windows with MgF-overcoated (for K-band) water-free silica (for GRACE, only external window needed) [TG, PIMS 269].
  14. Get rid of fold flat, leave DM hanging over edge of bench, removes 2 surfaces, and allows better interf view of DM.

NAOMI mechs

  1. Swap NAOMI simplex and osca-pre-stop positions
  2. WFS encoder track should be cleaned regularly (decided at AO meeting 1/06). Done yet?
  3. Left-side cover (as viewed from CCD end) for WFS needs refitting (after modifying, if necessary), to prevent dirt falling on track
  4. Y movement of probe stuck twice with 'soft limit' 10/03 Roy reported sticking, when y = 42, on moving to x = -37 (2/04).
  5. Occasional wandering/sticking of WFS focus movement. Wandering is particularly pernicious, observer doesn't notice, initially. Can stop by disabling encoders?

  6. Tune telescope oil dampers using accelerometer [EMM 4/03].

OASIS optics/mechs

  1. The 33-mm enlarger is vignetted by the mask (perforated plate) after the lenslet array. This enlarger is optimally focussed, but is delivering the wrong scale at the lenslet array. The field lens is probably a compromise for the 3 different enlargers used to date. A new enlarger is needed. Training with zemax model needed.

  2. Duck errors, particularly (all occurences logged): Grism wheel stuck again 1/3/05, required 11 attempts to move it. Wheels for which datum switches replaced by Juerg: grism 16/8/05. These duck errors appear not be a major problem, as of 1/06.

INGRID optics/mechs

See above for NAOMI-related INGRID issues e.g. throughput, vignetting by components before INGRID.

Test reports are in /data/instruments/INGRID/tests, filters, lin etc.

  1. High emissivity in K investigated 10/03-5/04, using new baffle, but no improvement. Comparisons confused by changes in ambient temperature. Bigger foreoptics might help, to reduce emission from mount, but then more ambient radiation enters. Ideal baffle gold on detector side, black on other side.
  2. Unreliable wheel movements, due to problems with motors, bearings, switches, addressed during 10/03-5/04 standdown. Phytron motors not used, being returned. Pupil-imager and focus-drive motors not replaced (yet).
  3. Filters - HBarr, KBarr removed as spares 4/04, replaced with blanks. Z replaced with blank. UCM replaced with Z. Barr J AO received 1/04, mfb to scan it at IAC. Cass K pupil stop changed with 2-hole pupil mask, Cass H with blank.
  4. New pupil stop IR fused silica + evaporated mask for central obstruction, to get rid of vanes. Substrate recieved from Crystran. Diameter of central obstr to be between 1.76 (naomi-obs) and 1.38 (osca-obs) (AZ email 2/1/04). Cost e1000
  5. Poor image quality at edge of field to be investigated using moving pinhole (astros). See earlier for details.
  6. Image PSF is insensitive to INGRID focus, but need to change focus on switching to pupil imager?
  7. Foreoptics being recoated 04/05 to reduce ghosting. Will also reduce emissivity a few % (i.e. much less than the 20% emissivity measured from foreoptics). Cass foreoptics were coated, NAOMI's weren't. Coating is hazardous, so plan is to coat and buy new set at same time, but costs e13000. Foreoptics were constructed for design f/ratio, or for NAOMI actual f/ratio?
  8. Need for automatic display and analysis (FWHM, FHHE, strehl of brightest star) of images as soon as they read out (crucial for 6+7/04).
  9. Grid of pinholes in front of INGRID needed to monitor image quality?

  10. Why is read noise worse than LIRIS? Gain still 6?
  11. Detector tilt is dominant contrib to poor image quality at Cass, but mfb advises won't be a problem with NAOMI (in accordance with experience). Hence not addressed 10/03 - 4/04 (orig to be done by SPW). Also (?) rotated 1 deg, not aligned N-E. Not a problem?
  12. Focus mask in pupil wheel (?) needs closer holes.
  13. 3/04 foreoptics lenses have moved slightly rel to one another, perp to axis. Probably not a serious problem.
  14. Better temperature sensor (Lakeshore diode) needed?
  15. New scripts to define appropriate pupil stops for each filter, when used with NAOMI. Most of narrow-band filters need to be used with broad-band blocking filters, esp Jcont.
  16. Hot pixels - apparent increase Jul 2003 because of clocking array while cooling.
  17. Data-reduction pipeline (no interface for INGRID or PFIP as of 12/03).
  18. Red pixel indicates counts > 28000, maybe we should have 2 colours e.g. yellow for > 12000.
  19. All current filters are as specified. HeI and Pagamma orders cancelled, because HeI low throughput, Pagamma long timescale. Z cancelled. Brgamma is OK after all.
  20. Web pages need advice on observing techniques (e.g. temp instability means need to take darks and flats at same time), list of published papers.
  21. Photograph INGRID components when open.
  22. Bright leak top-left corner?

OSCA optics/mechs

DIMM/MASS

Electronics/CCDs

  1. EPICS crate needs restoring (processor borrowed for naomisdsu). (SGR email 2/3/05)
  2. Cordless phone in GRACE.
  3. Repair dead segment 45, died 10.03. Also noted oscillating 8/03 by SELS, but fixed itself. Cables/cards need checking, reseating, impedance needs checking. (NB segment 47 was failing 9/01 due to bent pin on connector)
  4. Upgrade WFS to L3 'zero-noise' CCD.

  5. Better physical and electromagnetic (Faraday) protection for the DM cables; there was frequent pickup noise during the Jan 00 run. SCAM may have done elec-noise survey in GHRIL, which may indicate what kind of slding needed. Test by having strain-gauge rms vs time plot for a pass-thru segment up while trying different types and positions of shield.
  6. Look at earthing (see emails 22+31/7/03 from Andy Hide). [OPS team, after consultation with TG]
Computing
  1. Power-PC (Sunblade?) upgrade for NAOMI (3133), saves 12k/year in maintenance, sware effort to be contracted out?, cost 65k (at 4/03) [CSG, CFG]. Not yet funded 8/03. Craige 24/7/03 noted that navis sometimes has less than 60 MBytes (out of 1500) of virtual memory left. Lots of processes running, so e.g. netscape shouldn't be run on navis.
  2. Get fisba backed up (manually if necessary) [SJG/Don, PIMS 206]
  3. Upgrade 68k processors to power PCs to bring omc, sdsu etc. epics versions up to date (work by OSL now funded 9/03). Need to do before replacing navis? [SJG]
  4. Investigate (by start fy 2004) replacement of C40s by C67s, for higher loop speed [SJG].

Air-conditioning

  1. Humidity control to prevent rain in GRACE (as ~ 18/10/04). RH or dewpoint meassurement should trigger air-cond off, dehumidifier on?
  2. Cryostat-fill without opening double doors in-hand (KMD 21/10/04), auto-filling planned.

  3. Control of air cond (e.g. to set to 20%) + display of temp in control room (ordered by Doug ~ 20/10/03). Temperature logging also needed.
  4. Chiller unit outside GRACE to be replaced (no more work required in GRACE to complete air-cond) - couple of weeks standdown of GRACE needed for this (no cooling).
  5. Remove heat from acq and simplexing cameras.
  6. Mesh screens to control turbulence.
  7. Use INGRID with SDSU1 (noisy, 20e- rms, but no fan, can't be used at Cass?) or SDSU2 (as used 4/03, but vibrating fan, cooling plate won't fit). Remove fans not needed [AH].
  8. Cool GRACE to 10 degC all year round?

Documentation / spares / maintenance

  1. Order spares [TG]. First of all we need a spares list [OM, JR, Jim Baker] Secure storage needed for spares.
  2. Complete NAOMI change document, document alignment procedures [SJG].
  3. Ensure all documentation delivered by ATC and Durham. Highest priority is to create an html index of documents with current status of each (e.g. format, likely date of completion) so we can see where gaps are. Ideally, the titles in this index should become html links (and pdf as well, where appropriate) as the documents are released online. The index should indicate the date of creation of each document. [SJG] As of 6/03, change document is incomplete, alignment documentation doesn't exist yet. 1/06 - Olivier identified serial numbers of WFS-chassis drawing we appear not to have.

  4. Sort out (and try) maintenance procedures e.g. how do we clean the fragile pickoff mirror? Establish procedures (and frequency) for cleaning each surface. See AJL email 15/1/01. Reflectivities measured ~ 2002, holding up very well. Need to measure ~ once per year.

Post integration of GLAS

Checks of the AO system to carry out after integration of GLAS, to check that there has been no negative impact on the existing system:
Software checks
---------------
We could check most aspects of the systems by working through the standard
checks recipe on:
  http://www.ing.iac.es/~crb/wht/aorecipe1.html

In particular, the following should be checked:

- Check that the WFS reads out as expected in each of the currently-used
  modes.

- Check that the mirror can be laser-flattened, white-light flattened,
  and simplexed.  This will implicitly test data acquisition from the
  fisba, WFS and simplex cameras, control over the SG-loop/DM and over
  mechanisms in the WFS and USP.

- Record offsets on the WFS.

- Close the AO loop on a pinhole image, shift the NCU by 0.1 mm,
  check that the AO system compensates.

- Check that different reconstructors can be loaded.

- Check that the less-frequently used tools and diagnostic tools
  still run OK: e.g. AutoRectifyFlat, PrintTemperatures,
  EstimateFrameInterval.

- On-sky: close the loop, check that the correction achieved is
  reasonable and that the FSM traces (commanded movement vs time)
  look OK, check that autoguiding (signal to TCS) works.


Other checks
------------
- Stray light: compare light levels in broad-band imaging and
  spectroscopic modes on OASIS, check that counts have not
  increased since 1/06 measurements.  Painted-over LEDs may look
  black to the eye, but not to OASIS.

- Throughput: compare the throughput on the science arm with and
  without the GLAS beam-splitter in place, as a function of wavelength,
  using OASIS and INGRID.  Is it as expected?
  Check that there's nothing loose on the bench which could blow
  into the light path.

- Optical aberrations: check by comparison of simplexed flats
  that the beamsplitter isn't introducing any large changes
  in the wavefront.

- Flexure / accidental movement of optical components - does the
  pinhole still image on the WFS, OASIS, INGRID and AG7 in its
  usual position, and in focus at the standard filter-focus setting?

- Electronic noise: check for any increase in read noise or banding
  (or sensitivity to cables being moved on the bench) on each of
  the WFS, AG3, AG7, OASIS and INGRID detectors.

- Use the fisba to check for any change in vibration level with
  the GLAS components powered on and off.

- Check for excessive heat output from any of the new components.

- Check that the new components on the bench post no hazards to personnel
  (sharp edges and corners, trip hazards, exposed high voltage)
  or to other equipment (possible snagging of cables by moving parts,
  e.g. NCU).

Abandoned

  1. Fix the DM x stage, not worth doing?
  2. Field lens in front of OASIS. 10/03: not needed, vignetting was caused by misalign of lenslet array and adjacent baffle
  3. Access control for GRACE.
  4. Installation of OSCA pre-NAOMI 1" stop = substrate with 1" dichroic on top (reflecting IR). 2/3/04 P Doel advised no progress with Barr. Backup may be to produce thin gold beam splitter stop.
  5. Hinged protective cover over OSCA, and cable trunk moved aside to allow dichroic to be installed with less risk of dropping it. Soon according to KMD 21/10/04. Superceded by dichroic changer 11/05.
  6. Remote-alignment tool for OSCA

Done

This record (incomplete) started 2002 Jul 10 at the suggestion of SJG.

  1. TIP access to aocontrol1 via portserver [DBA].
  2. Software to generate star pairs of specified mag and separation for PSF calibration.
  3. Upgrade solaris (7 to 8?), after identifying any software dependent on old solaris [CFG/NAD, done 9/02].
  4. Command-line control of pickoff x,y [NAD, done 9/02].
  5. Put all software under CVS [SJG].
  6. Get netscape configured so that it can read .pdf files. [done 3/02]
  7. hhmmss names for mirror flats instead of hhmm (currently you have to wait one minute before saving another flat!) [NAD, PIMS 251]. Done 4/03.
  8. Option to display segment numbers on WFS display [NAD, PIMS 252]. Done 4/03.
  9. Button to save WFS image to FITS file [SJG/NAD, PIMS 261]. Done 4/03, but only works in mode 1.
  10. Display on gui the true loop speed for current wfs exposure time [SJG/NAD, PIMS 241]. Done 4/03.
  11. Stop topgui crashing if change mask in OSCA [SJG/NAD, PIMS 246]. Fixed 4/03.
  12. Explore why TopGui crashes so easily. NAD explored this 4/03, one cause is switching SDSU mode from outside sequencer, this now fixed. [SJG, PIMS 243].
  13. Display of DM temperature on top gui [NAD, PIMS 249].
  14. NCU lamp may need to provide more light for 4/03 run. Done.
  15. Commission SLODAR in GHRIL, with optical derotator. Done Apr 13-17 by Richard Wilson and Chris Saunter.
  16. Measure observing overheads accurately. Done 2002/3.
  17. Small GRACE things to be sorted out: no phone in optics area (fixed 5/03), access only possible at some azimuths, squeaky doors.
  18. What is actual size of INGRID pupil stops, correct for NAOMI? Reviewed by AZ SE RO IKS 8/03.
  19. Test new pointing model for GRACE, rms should be ~ 1.5 arcsec, to allow acquisition directly onto WFS. Done 7/03.
  20. Determine scale, field size, mag limit, orient of acquisition TV (3 arcmin field). Done 7/03 [CRB].
  21. Check illumination of pupil at old optical science port - what is origin of uneven illumination seen by SLODAR. Resolved mid-03 - beam was coming at angle 20 deg to that expected.
  22. Make sure INGRID exposure times are now understood [AZ].
  23. Investigate why Strehl of pinhole image drops on closing loop. How does this depend on the S:N of the offsets? What is minimum S:N on spots during closed-loop operation that doesn't reduce strehl? Found mid 03 to be due to Simplex positioning peak symmetrically to maximise 4 central pixels. Any small displacement decreases peak.
  24. Check best pinhole PSF thru current Z filter - diffraction-limited, or degraded by filter (i.e. is a new AO-quality Z filter needed?) [CRB] Done mid-03, no evidence that Z degrades wavefront.
  25. Identify cause of slow grab (0.7 Hz, instead of 2.2 Hz) during many nights in Aug [CB to test, ~ 3/03, PIMS 226]. CB's new simplex runs at 10 Hz, 8/03.
  26. Craige's dither script to work at non-zero PA. Done 4/03.
  27. Check vignetting of WFS caused by first doublet en route to OASIS (vignettes because focus position not quite where expected). Done 5/03, no vignetting.
  28. Move fisba control screen to WHT control room [SJG, PIMS 215]. Done 8/03.
  29. Topgui should display strain-gauge gain value [NAD]. Done early 03.
  30. Storage needed for dichroics [TG]. Done summer 03.
  31. Wrong gains for p1 p2 p3 and modes 11 12 corrected 8/03 [RM].
  32. Check can get 1250-microsec int time on WFS by raising decimate, test can be open loop [CRB, PIMS 240]. OK, done 8/03.
  33. 'Load flat on open' should be on main page [NAD, PIMS 248]. OK 8/03
  34. When the pickoff mirror x position is exactly 5, i.e. 5.00, it's not possible to change the y position! [CRB to check if still a problem, PIMS 253]. OK 8/03.
  35. Rotate MIT3 cryostat 90 deg, so to optimise direction of charge smearing. Done 9/03. MIT hold time checked, ~ 14 hours.
  36. Identify the source of the 0.2 sec overhead in the grab (not seen when grabbing direct from crater. If transfer protocol is slowing the grab investigate possibility of calculating merit function in ultradas and sending DM instructions direct from udas to aocontrol1 [SJG (maybe RAB after 3/03), PIMS 244]. Craige's new 10-Hz simplex tested. Done.
  37. Measure best PSF on MIT/LL-CCD image of pinhole (and compare with that measured at INGRID). 2 pixels achieved 10/03 on each of INGRID, SBIG, MIT3
  38. Derot elements now arrived, 8/03, to be installed 10/03.
  39. Compare, after good simplex, peak counts H vs HBarr, Ks vs KsBarr. Done for both 10/03, no change (see web page for results).
  40. Indication of when DM segments and FSM hit limits. Better a new display with TCL/TK [CB, PIMS 225]. There is an indication if FSM hits limit, but not realtime. Done by Craige, ~ 8/03.
  41. Rotate MIT CCD by 90 deg. Done 10/03.
  42. Change navis display to 2-monitor setup [CRB/Don, PIMS 242]. 2 screens now available, summer 03.
  43. Why does 50:50 dichroic deliver single image of pinhole, but double image of pupil (noted by Roy). In fact, there is a double image of pinhole too, second being out of focus 2", and 15" left on INGRID, refln from back surface dichroic? (crb)
  44. Mode9 button on launcher [SJG/CB, PIMS 257]. There's now a button on the config menu (summer 03).
  45. Check can close loop on M32 with poor-seeing lenslet (done 7+8/04).
  46. Test vignetting by OSCA optics by imaging with OASIS on-sky, done 9/04.
  47. Meaure FSM frequency-response performance, which may be limiting NAOMI performance, and/or rejection of telescope oscillations above given amplitude or frequency [SJG]. See AJL email 15/10/03. TG suggests 10/03 voice-coil-driven transp plate in front of pinhole, amplitide ~ +- 1". Status 9/04: poor response traced to 1-msec delay introduced by frequency filter (added after delivery from Zeiss), now fixed.
  48. Determine new value of focus-corrector value required for each filter. 10/03: The pinhole image is in focus on the SBIG camera when the focus corrector (relay lenses) is set to 4.05. The pinhole viewed through the OASIS MR591 filter is in focus on MIT3 when the focus corrector is set to 7.0 rather than the former value (and default on gui) of 5.91. It is assumed that this offset of 1.1 applies to all filters. Unfortunately, this means that the correct focus for some filters is at values > 8.5, the limit of the drive. We need to move the focus corrector by about 3 mm, so that the focus values are in the middle of the range 0 - 8.5 mm, rather than close to 8.5 mm. The default values on the gui are apparently set to wavel(A)/1000, i.e. wrong.
  49. Commission AG5 simplexing camera, currently (3/04) awaiting new bracket (prob ready by 6/04). Will allow accurate focussing. Done ~ 8/04.
  50. Diamond pattern-noise on WFS CCD 9/03, possibly due to CCD connection being strained when camera stage is moved DR 15972 . Fixed sumer 04, controller was overheating.
  51. High bias level on AG5 affecting simplex result. Bias reduced from 6100 to 800 by SMT 8/04.
  52. Make AutoWhite use current exposure time (instead of changing to 200 msec). Done.
  53. Binning on MIT3 CCD (done 11/3/05 SMT).
  54. XOASIS package to be available at ING. [Done ~mid 04]
  55. Correct WFS display for TT mode. [Done 11/04, NAD]
  56. Modify Craige's guide-star acquisition tool so that can acquire guide star on WFS, then offset (telescope and pickoff) so that science object is centred on detector [SE/CB, PIMS 105].
  57. Automatic cropping of simplexes from AG2 and other CCDs to remove extra bright columns, which may affect the calculated merit function [Jure 1/05].
  58. Stop FISBA ftp logs filling up the navis /var disk (e.g. switch from ftp to nfs, see Paul Clark's email 25/6/01, or use SAMBA). Filled up 3 times so far during Feb 2003. [CB, PIMS 217] [OK 5/05]
  59. Restore f77 licence on navis, expires monthly because only demo licence. [Done ~ 3/05]
  60. Stage lamps on top-end ring, for OASIS dome flats (emailed KMD 24/8/04, see also RMcD email 21/9/04). In hand at 21/10/04 (KMD). [Done early 05]
  61. Check calibration of WHT internal and mirror thermometers, which currently imply mirror seeing on 20% nights (req KMD 25/9/04). In-hand (Berto) 21/10/04 acc to KMD. [Done early 05].
  62. Make fisba z gain default to 0. [Done 5/05 by Jure]
  63. Warning from fisba gui to restore airco to 100% after laser-flattening. [Done 5/05 by Jure]
  64. Identify best focus position to use in INGRID. Various combinations of tel and INGRID focus yield in-focus image. Need a combination which gives good quality images, plate scale 0.238-0.240 "/pixel. 260 mic of INGRID focus = 0.1 mm of tel focus.
  65. Commission dichroic changer, when installed (~ 9/05, acc to mvdh email 28/10/04, for test scripts see spico email 5/11/04). Done 11/05.
  66. Convert NAOMI observing recipes into manual for local staff [CRB]. Done early 05.
  67. OASIS service form needed, reminder sent 20/10/03 + 20/4/04 + 9/04 [PLeisy]. Done mid 05.
  68. Specify instrument-control gui [CRB]. Done 05?
  69. Training for ING engineers. Done early 05.
  70. aocontrol1 replacement, will make loop faster? [CSG, CFG, Durham] Done 04, 05?
  71. 33-mm wouldn't focus 26/4/05 mfb found spacers to be wrong, has removed them, also found lens to be wrong way round ~ mid 05.
  72. Align dichroic 4 with others [MFB/TG]. All 10 supposed to be co-aligned 10/04, but Roy 2/10/03 found pupil movement 1 segment on changing between 50:50 and IR dichroics. IR dichroic still in old mount, needs to be in same type of mount as the others. Realignment to be done after delivery of cone/v/flat mount, not yet ordered 21/01/04 acc mfb. Realigned for use in dichr changer aut 05.
  73. Dichroic changer, and fallback to manual operation. Done aut 05 (but no fallback to manual).
  74. Some WFS exposure times are unavailable with certain modes. Can menu easily be modified to reflect this dependence? Alternatively a warning message if one attempts to select an inappropriate exposure time [SJG/NAD, PIMS 259].
  75. Lyot stop doesn't track rotn of sec vanes with correct direction, fixed 9/05.
  76. Basketball under flat has been removed. Probably not needed at present (RO email 3/10/05).
  77. After dichroic changer installed, OSCA wouldn't fit. OSCA detent mech modified by MVDH 1/06.
  78. AutoWhite should log results to a file, and peakH etc should be stored in specific directory, not current, and not cdflat, done (Jure late 05)
  79. Option to set WFS exposure time manually, so can set with better precision. [CB to investigate restrictions, CRB to decide what new exposure times would be most useful, PIMS 232]. Done (spico?) late 05.
  80. OASIS MR454 filter flopping in mount, needs packing, advised mfb 16/6/05. Fixed by mfb 7/05.
  81. LIRIS bright glow needs covering up (req KMD 21/1/05). Done 05.
NAOMI technical information · WHT · Astronomy · ING

ING Logo

Last Updated: 2006 Dec 10
Chris Benn crb@ing.iac.es